# Model Tampering Attacks Enable More Rigorous Evaluations of LLM Capabilities

### v2

@@ Abstract
Evaluations of large language model (LLM) risks and capabilities are increasingly being incorporated into AI risk management and governance frameworks. Currently, most risk evaluations are conducted by designing inputs that elicit harmful behaviors from the system. However, this approach suffers from two limitations. First, input-output evaluations cannot evaluate realistic risks from open-weight models. Second, the behaviors identified during any particular input-output evaluation can only lower-bound the model’s worst-possible-case input-output behavior. As a complementary method for eliciting harmful behaviors, we propose evaluating LLMs with model tampering attacks which allow for modifications to latent activations or weights. We pit state-of-the-art techniques for removing harmful LLM capabilities against a suite of 5 input-space and 6 model tampering attacks. In addition to benchmarking these methods against each other, we show that (1) model resilience to capability elicitation attacks lies on a low-dimensional robustness subspace; (2) the attack success rate of model tampering attacks can empirically predict and offer conservative estimates for the success of held-out input-space attacks; and (3) state-of-the-art unlearning methods can easily be undone within 16 steps of fine-tuning. Together these results highlight the difficulty of suppressing harmful LLM capabilities and show that model tampering attacks enable substantially more rigorous evaluations than input-space attacks alone. [^1] We release 64 models at https://huggingface.co/LLM-GAT.
@@ 2 Related Work
Capability elicitation and evaluation: Research on adversarial capability elicitation (Hofstätter et al., 2025) in LLMs has primarily been done in the context of machine unlearning (Liu et al., 2024a; Barez et al., 2025) and jailbreaking (Yi et al., 2024b). Here, we experiment in these two domains. However, capability elicitation has also been researched in the context of backdoors/trojans (Zhao et al., 2024), “password-locked models” (Greenblatt et al., 2024; Hofstätter et al., 2025), and “sandbagging” (van der Weij et al., 2024). In the unlearning field, several recent works have used adversarial methods to evaluate the robustness of unlearning algorithms (Patil et al., 2023; Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Liu et al., 2024a; Zhang et al., 2024; Liu et al., 2024b). Here, we build off of Li et al. (2024b) who introduce WMDP-Bio, a benchmark for unlearning dual-use biotechnology knowledge from LLMs. In the jailbreaking field, many techniques have been developed to make LLMs comply with harmful requests (Shayegani et al., 2023; Yi et al., 2024b; Jin et al., 2024; Chowdhury et al., 2024; Lin et al., 2024). Here, we experiment with 9 open-source LLMs and a set of gradient-guided, perplexity-guided, and prosaic techniques from the adversarial attack literature (see Table 1).
@@ 3 Methods
PullBack & proJect (PB&J) Anonymous (2025)
@@
Table 1: Table of capability elicitation (attack) and capability suppression (defense) defense. We consider defenses in two different settings: (top) unlearning approaches that remove hazardous bio-knowledge and (bottom) refusal-tuned models that resist jailbreaks.
@@
Our approach: pitting capability suppression defenses against capability elicitation attacks. For unlearning experiments, we experiment with 65 models trained using 8 different unlearning methods. For jailbreaking experiments, we experiment with 9 models off the shelf from prior works. In both cases, we pit these defenses against a set of 11 input-space and model tampering attacks to either elicit ‘unlearned’ knowledge or jailbreak the model. In Table 1, we list all unlearning methods, off-the-shelf models, and attacks we use. Since the input-space attacks that we use are held out, we treat them as proxies for novel input-space attacks in our evaluations (see also Hofstätter et al. (2025)).
@@
Defenses: machine unlearning methods: We unlearn dual-use bio-hazardous knowledge on Llama-3-8B-Instruct Dubey et al. (2024) with the unlearning methods listed in Table 1 and outlined in Section A.2.1. For all methods, we train on 1,600 examples of max length 512 from the bio-remove-split of the WMDP ‘forget set’ (Li et al., 2024b), and up to 1,600 examples of max length 512 from Wikitext as the ‘retain set’. For the 8 unlearning methods listed in Table 1, we take 8 checkpoints evenly spaced across training. Finally, we also use the public release of the “TAR-v2” model from Tamirisa et al. (2024) as a 9th TAR model. In total, the 8 checkpoints each from the 8 methods we implemented plus the TAR model from Tamirisa et al. (2024) resulted in 65 models.
@@
Defenses: refusal fine-tuned models: For jailbreaking experiments, we use the 9 fine-tuned Llama3-8B-Instruct models off the shelf listed in Table 1. The first 8 are all fine-tuned for robust refusal of harmful requests. Of these, ‘RR’ (Zou et al., 2023a) and ‘LAT’ (Sheshadri et al., 2024) are state-of-the-art for open-weight jailbreak robust models (Li et al., 2024a; Haize Labs, 2023). The final ‘Orenguteng’ model was fine-tuned to be ‘helpful-only’ and comply even with harmful requests. We discuss these models in more detail in Section A.3.
@@
Attacks: capability elicitation methods: We use 5 input-space attacks and 6 model tampering attacks on our unlearned models. We use these attacks (single-turn) to increase dual-use bio knowledge (as measured by WMDP-Bio performance (Li et al., 2024b)) for unlearning experiments and to elicit compliance with harmful requests (as measured by the StrongReject AutoGrader (Souly et al., 2024)) for jailbreaking experiments. We selected attacks based on algorithmic diversity and prominence in the state of the art. We list all 11 attacks in Table 1. In all experiments, we produce universal adversarial attacks optimized to work for any prompt. This allows us to attribute attack success to capability elicitation rather than answer-forcing from the model (e.g., Fort (2023)). For descriptions and implementation details for each attack method, see Appendix A. Finally, we also used two proprietary attacks – one for unlearning experiments and one for jailbreaking experiments which we will describe in Section 4.
Attacks: data
@@
Attacks: model tampering attacks are efficient. In Table 3, we show the number of forward and backward passes used in our implementations of attacks. Model tampering attacks were more efficient than state-of-the-art input-space attacks.
@@ 4.1.2 Model robustness exists on a low-dimensional subspace
We perform PCA, weighting models by their unlearning score. First, to understand the extent to which some attacks offer information about others, we analyze the geometry of attack successes across our 65 models. We perform weighted principal component analysis on the WMDP-Bio improvements achieved by all 11 attacks on all 65 model checkpoints. We first constructed a matrix $A$ with one row per model and one column per attack. Each $A_{ij}$ corresponds to the increase in WMDP-Bio performance in model $i$ under attack $j$. We then centered the data and multiplied each row $A_{i}$ by the square root of the unlearning score: $\sqrt{S_{\text{unlearn}}(A_{i})}$. This allowed for models to influence the analysis in proportion to their unlearning score.
@@ 5 Discussion
  • Can models be robust to tampering attacks? This paper and concurrent work (Qi et al., 2024b) show that even defenses designed to make models robust to tampering can be easily undone. We are currently working to better understand tampering robustness and improve the extent to which models can be made robust to tampering attacks.
  • What mechanisms underlie robust capability removal? We are interested in future work to mechanistically characterize weak vs. robust capability suppression. We hope that the 64 models we release help to lay the groundwork for this.
@@ A.2.1 Unlearning Methods
  • PullBack & proJect (PB&J): Anonymous (2025) is an unlearning algorithm which learns a set of projections which on activations space which maximally harm performance on the the forget set while minimally perturbing model outputs on a broad retain distribution.
To adhere to the implementations from the works introducing each method, we use full fine-tuning (not LoRA) for RMU, RMU-LAT, RepNoise, TAR, and PB&J, and LoRA for GradDiff, ELM, RR.
@@ A.4 Attack Methods and Implementation
GCG (Zou et al., 2023b) performs token-level substitutions to an initial prompt by evaluating the gradient with respect to a one-hot vector of the current token. We implemented both time-bounded attacks on each unlearned model and transfer attacks using prefixes from one model to attack others. Unless otherwise specified, we report the mean performance of each gradient-guided attack.
@@
Like GCG, AutoPrompt (Shin et al., 2020) performs a gradient-guided search over input tokens to design universal adversarial prompts.
@@ Appendix D Full Jailbreaking Results
Figure 15: All safety-tuned models could be successfully jailbroken, particularly by fine-tuning and Cascade attacks. We evaluate safety-tuning methods and jailbreak attacks. Left: The ‘Baseline’ measures the compliance rate to direct harmful requests. Right: Increase in harmful response rate after attack. All safety-tuning methods were vulnerable to elicitation of suppressed capabilities, especially by finetuning and Cascade attacks (rightmost columns).

### v3

@@ Abstract
Evaluations of large language model (LLM) risks and capabilities are increasingly being incorporated into AI risk management and governance frameworks. Currently, most risk evaluations are conducted by designing inputs that elicit harmful behaviors from the system. However, this approach suffers from two limitations. First, input-output evaluations cannot fully evaluate realistic risks from open-weight models. Second, the behaviors identified during any particular input-output evaluation can only lower-bound the model’s worst-possible-case input-output behavior. As a complementary method for eliciting harmful behaviors, we propose evaluating LLMs with model tampering attacks which allow for modifications to latent activations or weights. We pit state-of-the-art techniques for removing harmful LLM capabilities against a suite of 5 input-space and 6 model tampering attacks. In addition to benchmarking these methods against each other, we show that (1) model resilience to capability elicitation attacks lies on a low-dimensional robustness subspace; (2) the success rate of model tampering attacks can empirically predict and offer conservative estimates for the success of held-out input-space attacks; and (3) state-of-the-art unlearning methods can easily be undone within 16 steps of fine-tuning. Together these results highlight the difficulty of suppressing harmful LLM capabilities and show that model tampering attacks enable substantially more rigorous evaluations than input-space attacks alone. [^1] We release models at https://huggingface.co/LLM-GAT.
@@ 2 Related Work
Capability elicitation and evaluation: LLMs are commonly developed by simply training them to behave desirably (e.g., with RLHF (Casper et al., 2023)), but in this paper, we focus on testing targeted defenses against known, harmful behaviors. Research on adversarial capability elicitation (Hofstätter et al., 2025) in LLMs has primarily been done in the context of machine unlearning (Liu et al., 2024a; Barez et al., 2025) and jailbreaking (Yi et al., 2024b). Here, we experiment in these two domains. However, capability elicitation has also been researched in the context of backdoors/trojans (Zhao et al., 2024), “password-locked models” (Greenblatt et al., 2024; Hofstätter et al., 2025), and “sandbagging” (van der Weij et al., 2024). In the unlearning field, several recent works have used adversarial methods to evaluate the robustness of safety fine-tuning and unlearning algorithms (Patil et al., 2023; Lynch et al., 2024; Łucki et al., 2024; Hu et al., 2024; Liu et al., 2024a; Zhang et al., 2024; Liu et al., 2024b; Wei et al., 2024). Here, we build on Li et al. (2024b) who introduce WMDP-Bio, a benchmark for unlearning dual-use biotechnology knowledge from LLMs. In the jailbreaking field, many techniques have been developed to make LLMs comply with harmful requests (Shayegani et al., 2023; Yi et al., 2024b; Jin et al., 2024; Chowdhury et al., 2024; Lin et al., 2024). Here, we experiment with 9 open-source LLMs and a set of gradient-guided, perplexity-guided, and prosaic techniques from the adversarial attack literature (see Table 1).
@@ 3 Methods
K-FAC for Distribution Erasure (K-FADE) McKinney et al.
@@
Table 1: Table of capability elicitation (attack) and capability suppression (defense) methods. We consider defenses in two different settings: (top) unlearning approaches that remove hazardous bio-knowledge and (bottom) refusal-tuned models that resist jailbreaks.
@@
Our approach – pitting capability suppression defenses against capability elicitation attacks. Here, we study capability suppression methods that depend on both removing knowledge from the model (unlearning) and teaching the model to robustly refuse (jailbreaking) requests. For unlearning experiments, we experiment with 65 models trained using 8 different unlearning methods. For jailbreaking experiments, we experiment with 9 models off the shelf from prior works. In both cases, we pit these defenses against a set of 11 input-space and model tampering attacks to either elicit ‘unlearned’ knowledge or jailbreak the model. In Table 1, we list all unlearning methods, off-the-shelf models, and attacks we use. Since the input-space attacks that we use are held out, we treat them as proxies for novel input-space attacks in our evaluations (see also Hofstätter et al. (2025)).
@@
Defenses – machine unlearning methods: We unlearn dual-use bio-hazardous knowledge on Llama-3-8B-Instruct Dubey et al. (2024) with the unlearning methods listed in Table 1 and outlined in Section A.2.1. For all methods, we train on 1,600 examples of max length 512 from the bio-remove-split of the WMDP ‘forget set’ (Li et al., 2024b), and up to 1,600 examples of max length 512 from Wikitext as the ‘retain set’. For the 8 unlearning methods listed in Table 1, we take 8 checkpoints evenly spaced across training. Finally, we also use the public release of the “TAR-v2” model from Tamirisa et al. (2024) as a 9th TAR model. In total, the 8 checkpoints each from the 8 methods we implemented plus the TAR model from Tamirisa et al. (2024) resulted in 65 models.
@@
Defenses – refusal fine-tuned models: For jailbreaking experiments, we use the 9 fine-tuned Llama3-8B-Instruct models off the shelf listed in Table 1. The first 8 are all fine-tuned for robust refusal of harmful requests. Of these, ‘RR’ (Zou et al., 2023a) and ‘LAT’ (Sheshadri et al., 2024) are state-of-the-art for open-weight jailbreak robust models (Li et al., 2024a; Haize Labs, 2023). The final ‘Orenguteng’ model was fine-tuned to be ‘helpful-only’ and comply even with harmful requests. We discuss these models in more detail in Section A.3.
@@
Attacks – capability elicitation methods: We use 5 input-space attacks and 6 model tampering attacks on our unlearned models. We use these attacks (single-turn) to increase dual-use bio knowledge (as measured by WMDP-Bio performance (Li et al., 2024b)) for unlearning experiments and to elicit compliance with harmful requests (as measured by the StrongReject AutoGrader (Souly et al., 2024)) for jailbreaking experiments. We selected attacks based on algorithmic diversity and prominence in the state of the art. We list all 11 attacks in Table 1. In all experiments, we produce universal adversarial attacks optimized to work for any prompt. This allows us to attribute attack success to capability elicitation rather than answer-forcing from the model (e.g., Fort (2023)). For descriptions and implementation details for each attack method, see Appendix A. Finally, we also used two proprietary attacks – one for unlearning experiments and one for jailbreaking experiments which we will describe in Section 4.
Attacks – data:
@@
Attacks – model tampering attacks are efficient. In Table 3, we show the number of forward and backward passes used in our implementations of attacks. Model tampering attacks were more efficient than state-of-the-art input-space attacks.
@@ 4.1.2 Model robustness exists on a low-dimensional subspace
We perform PCA, weighting models by their unlearning score. First, to understand the extent to which some attacks offer information about others, we analyze the geometry of attack successes across our 65 models. Previously Wei et al. (2024) found that a model’s vulnerability to pruning and low-rank modifications both relate with the brittleness of its safety fine-tuning. Here, we extend on this finding with more attacks and subspace analysis. We perform weighted principal component analysis on the WMDP-Bio improvements achieved by all 11 attacks on all 65 model checkpoints. We first constructed a matrix $A$ with one row per model and one column per attack. Each $A_{ij}$ corresponds to the increase in WMDP-Bio performance in model $i$ under attack $j$. We then centered the data and multiplied each row $A_{i}$ by the square root of the unlearning score: $\sqrt{S_{\text{unlearn}}(A_{i})}$. This allowed for models to influence the analysis in proportion to their unlearning score.
@@ 4.2 Jailbreaking Experiments
Figure 8: All safety-tuned models could be successfully jailbroken by fine-tuning and Cascade attacks. We evaluate safety-tuning methods and jailbreak attacks. Left: The ‘Baseline’ measures the compliance rate to direct harmful requests. Right: Increase in harmful response rate after attack. All safety-tuning methods were vulnerable to elicitation of suppressed capabilities, especially by finetuning and Cascade attacks (rightmost columns).
@@ 5 Discussion
  • Can models be robust to tampering attacks? This paper and concurrent work (Qi et al., 2024b) show that even defenses designed to make models robust to tampering can be easily undone. We are currently working to better understand tampering robustness and improve the extent to which models can be made robust to tampering attacks through pretraining interventions (e.g. Maini et al., 2025), knowledge corruption (e.g. Wang et al., 2025), and improvements in unlearning algorithms (e.g. Siddiqui et al., 2025).
  • What mechanisms underlie robust capability removal? We are interested in future work to mechanistically characterize weak vs. robust capability suppression. We briefly worked to test the hypothesis that the activation rank difference (across the forget set) between a base and unlearned model would correlate with unlearning robustness. However, we found this not to be the case, and leave further investigation to future work. We hope that the 64 models we release help to lay the groundwork for this.
@@ A.2.1 Unlearning Methods
  • K-FAC for Distribution Erasure (K-FADE): McKinney et al. is an unlearning algorithm which learns a set of projections which on activations space which maximally harm performance on the the forget set while minimally perturbing model outputs on a broad retain distribution.
To adhere to the implementations from the works introducing each method, we use full fine-tuning (not LoRA) for RMU, RMU-LAT, RepNoise, TAR, and K-FADE, and LoRA for GradDiff, ELM, RR.
@@ A.4 Attack Methods and Implementation
GCG (Zou et al., 2023b) performs token-level substitutions to an initial prompt by evaluating the gradient with respect to a one-hot vector of the current token. Unlike standard GCG, which is typically used to make a model output a specific string, we used a universal version of GCG, optimized over a set of examples to elicit a more general harmful behavior (e.g., giving correct responses to biology questions). We implemented both time-bounded attacks on each unlearned model and transfer attacks using prefixes from one model to attack others. Unless otherwise specified, we report the mean performance of each gradient-guided attack.
@@
Like GCG, AutoPrompt (Shin et al., 2020) performs a gradient-guided search over input tokens to design universal adversarial prompts. As with GCG, we create universal versions of these attacks using a set of examples.